Two rockets A and B depart from earth at a steady speed of 0.6c in opposite directions, having synchronized clocks with each other and with earth at departure. After one year as measured in Earth’s frame, rocket B emits a light signal. At what time, in the frames of earth and rockets A and B, does rocket A receive the signal? Work out the problem in all three frames and verify the answers are same. Use unit of distance as light year and that of time as year.


1
Expert's answer
2021-01-07T10:54:59-0500

Solution:


In this problem we have 3 inertial frames to consider, so we will use coordinates x and t to denote quantities measured in the Earth`s frame, "x_A", and "t_A" to denote quantities measured in A`s frame, and "x_B" and "t_B" to denote quantities measured in B`s frame. We will assume that, according to an observer in the Earth`s frame, rocket A is travelling at speed v in the positive x-direction and rocket B in the negative x-direction. With these assumptions, we can write down the Lorentz transformations:


"x=\gamma(x_A+vt_A)" , "x_A=\gamma(x-vt)"

"t=\gamma(t_A+\dfrac{vx_A}{c^2})," "t_A=\gamma(t+\dfrac{vx}{c^2}),"

"x=\gamma(x_B-vt_B)," "x_B=\gamma(x+vt),"

"t=\gamma(t_B-\dfrac{vx_B}{c^2}), \; t_B=\gamma(t+\dfrac{vx}{c^2}),"


where

"v=0.6c"


and

"\gamma=(1-\dfrac{v^2}{c^2})^{-\frac{1}{2}}=\dfrac{5}{4}"


It will be convenient to use a system of units in which time is measured in years and distance in light-years, in which case c has a value of 1. Again, it is helpful to draw a space-time diagram in the Earth`s frame of reference, as shown in the following picture:




so at time "t\gt1" the x-coordinate of the light signal is



"x = -0.6+(t-1)=t-1.6"

At time t, the x-coordinate of rocket A is



"x = +0.6t"


The x-coordinates must be equal at event 2, so for this event, we must have



"t-1.6=0.6t"


hence t=4. Thus according to an observer in the Earth frame, event 2 occurs after 4 years. The x-coordinate of this event us therefore



"0.6\cdot 4=2.4"


We can now substitute these coordinates into the Lorentz transformation formulae to find the time coordinates of event 2 in A`s and B`s frame of reference:



"t_A=\gamma(t-vx)"


(remember that c=1)


"=\dfrac{5}{4}(4-0.6\cdot2.4)=3.2"


so that, according to A, event 2 occurs after 3.2 years.



"t_B=\gamma(t+vx)=\dfrac{5}{4}(4+0.6\cdot2.4)=6.8"



so according to B, event 2 occurs after 6.8 years.